what is the percent by mass of each element in sucrose
1
Expert's answer
2013-04-10T08:36:16-0400
The molecular formula of sucrose is C12H22O11

1) The molar weight of sucrose is
MW = 12 ∙ MW(C) + 22 ∙ MW(H) + 11 ∙ MW(O)
MW(sucrose) = 12 ∙ 12 + 22 ∙ 1 + 11 ∙ 16 = 342 g/mol

2) The percent by mass is
ω% = m(element) / m (sucrose) ∙ 100%
ω%(C) = 12 ∙ 12 / 342 ∙ 100% = 42.11 %
ω%(H) = 22 ∙ 1 / 342 ∙ 100% = 6.43 %
ω%(O) = 11 ∙ 16 / 342 ∙ 100% = 51.46 %

Answer: ω%(C) = 42.11 %
ω%(H) = 6.43 %
ω%(O) = 51.46 %